I think i have the right answer, but the problem is that my book dosen't really explain well enough on how to write the equation.
The band leader wants to line up the 92 members in 4 rows so that each row has two more than the row before. How many members whould be in the shortest row?
I'm sure the answer is 17. But I can't figure out for the life of me how to write the equation. I've been working in it all day.

Answer by stanbon(75887)

The band leader wants to line up the 92 members in 4 rows so that each row has two more than the row before. How many members whould be in the shortest row?
-------------
Let the shortest line have x members
The next is x+2 members
Then x+2+2=x+4 members
And finally x+4+2 = x+6 members
------------------------
EQUATION;
x + x+2 + x+4 + x+6 = 92
4x+12 = 92
4x = 80
x = 20 (number in the shortest line)
==============
Cheers,
Stan H.